OpenStudy (anonymous):

34 = -13.6 + (3.4c) + 1.7c

OpenStudy (anonymous):

is it 9.3?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

5.1c = 34 + 13.6 = 47.6 ->c = 47.6 /5.1 = 9.3

OpenStudy (anonymous):

9.333333=9.33

OpenStudy (anonymous):

1c=9.33333333

OpenStudy (anonymous):

47.6=[3.4+1.7=5.1]=c c=9.3333

OpenStudy (anonymous):

THE ANSWER CHOICES ARE: A. –8 B. 2 C. –10 D. –4

OpenStudy (anonymous):

helppppppppppp

OpenStudy (anonymous):

please